OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) + 6. 1 4 5 25

jigglypuff314 (jigglypuff314):

to solve for x: 3(x + 1) = -2(x - 1) + 6 first, distribute so 3(x) + 3(1) = -2(x) + (-2)(-1) + 6 then simplify 3x + 3 = -2x + 2 + 6 3x + 3 = -2x + 8 then subtract 3 from both sides -3 -3 3x + 0 = -2x + 5 then add 2x to both sides +2x +2x 5x = 0 + 5 then divide both sides by 5 /5 /5 x = 5/5 then reduce and you have your answer x =